# On the Foundation of Distributionally Robust Reinforcement Learning

### v3

@@ 3.1.1 Deterministic Controller Policies
An important special case is when $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}=\left\{{\delta_{\left\{{a}\right\}}:a\in A}\right\}$. A deterministic rule is naturally uniquely identified with a function $d_{t}:\mathbf{H}_{t}\rightarrow A$. We will always assume such identification and use $d_{t}(h_{t})$ as a measure or an action interchangeably.
As clarified in Wiesemann et al. (2013), in the S-rectangular setting with $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}$ for $\Pi^{\mathrm{C}}$ and $\mathcal{Q}=\mathcal{P}(\mathcal{A})$ for $\Pi$, $v^{*}(\mu,\Pi^{\mathrm{C}},\mathrm{K})<v^{*}(\mu,\Pi,\mathrm{K})$ could happen for $\mathrm{K}=\mathrm{K}_{\mathrm{H}}^{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$. Thus, there could be no deterministic policy that is optimal for the controller that is allowed to take randomized actions. However, if one is to constrain the controller to take policies that chooses deterministic action, it is still instrumental to define the $q$-function and its Bellman equation as in the classical MDP settings.
@@
Define the DR Bellman equation for the $q$-function as
$q(s,a)=r(s,a)+\gamma\inf_{p_{s}\in\mathcal{P}_{s}}p_{s,a}\left[\max_{b\in A}q(\cdot,b)\right].$ (3.3)
It is important to note that for each $s$, the infimum chooses a $p_{s}=\left\{{p_{s,a}:a\in A}\right\}$ for each $a\in A$; i.e. such choice could be different across $a$. Effectively, for different controller action’s action $a\in A$,
$q(s,a)=r(s,a)+\gamma\inf_{p\in\left\{{p_{s,a}:p_{s}\in\mathcal{P}_{s}}\right\}}p\left[\max_{b\in A}q(\cdot,b)\right].$ (3.4)
In particular, $\left\{{p_{s,a}:p_{s}\in\mathcal{P}_{s}}\right\}$ is the marginalization (see (2.11)) of $\mathcal{P}_{s}$.
With the DR Bellman equation for the $q$-function defined, we explore the connection of its solution with that of the Bellman equation of the value function (3.1). Let $u^{*}$ denote the unique solution of (3.1). Then, we define the optimal $q$-function as follows: for each $(s,a)\in S\times A$
$q^{*}(s,a)=r(s,a)+\gamma\inf_{p_{s}\in\mathcal{P}_{s}}p_{s,a}[u^{*}].$
The following corollary links $q^{*}$ and $u^{*}$ to the solution of (3.4), and hence establishes an alternative equivalent formulation of the DPP in the case of deterministic controller.
Corollary 1.1.
$q^{*}$ is the unique solution to (3.4) and $u^{*}(\cdot)=\max_{a\in A}q^{*}(\cdot,a)$ satisfying Theorem 1.
Proof of Corollary 1.1.
The uniqueness follows from the same contraction argument as in the proof of Proposition 1. Moreover,
$u^{*}(s)$ $=\sup_{d\in\mathcal{Q}^{\mathrm{D}}}d[r(s,\cdot)]+\gamma\inf_{p_{s}\in\mathcal{P}_{s}}d\otimes p_{s}[u^{*}]$
$=\max_{a\in A}r(s,a)+\gamma\inf_{p_{s}\in\mathcal{P}_{s}}p_{s,a}[u^{*}]$
$=\max_{a\in A}q^{*}(s,a).$
Plug this into (3.3), one sees that $q^{*}$ is indeed the solution, hence equivalently solves (3.4). Lastly, Theorem 1 still holds in the context of $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}$. ∎
This corresponds to all presences of six check marks in Table 4. Subsequently, we will examine and establish the optimality attributed to the greedy policies derived from the optimal $q^{*}$-function in Remark 5. Moreover, we will provide counterexamples in Section 5.3, substantiating the red crosses within Table 4.
@@ 3.2 The Inf-Sup Equation and the Minimax Theorem
In this section, we first go back to the scenario where $\mathcal{Q}\subset\mathcal{P}(\mathcal{A})$ denotes an arbitrary subset containing action distributions. Our focus shifts towards assessing the validity of the DPP given by the maxmin Bellman equation (3.1). This investigation encompasses an additional assumption wherein the solution $u^{*}$ of (3.1) concurrently satisfies (3.2). Apart from establishing the DPP within Table 2, our pursuit involves addressing the subsequent query: Does it always hold at this level of generality that the optimal maxmin value in (2.16) conforms to a minimax theorem? In other words, does interchanging the sup and inf operations in (2.16) always preserve the value? The answer is yes, and is proved in Theorem 2. From this theorem, the validity of Table 2 is then established.
@@
Let $u^{*}$ be the solution of the DR Bellman equation (3.1). Assume $u^{*}$ further satisfies the inf-sup equation (3.2). Then, the following properties holds for all $\mu\in\mathcal{P}(\mathcal{S})$ and any mixture of $\mathrm{K}=\mathrm{K}_{\mathrm{H}}^{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ and $\Pi=\Pi^{\mathrm{C}}_{\mathrm{H}},\Pi^{\mathrm{C}}_{\mathrm{M}},\Pi^{\mathrm{C}}_{\mathrm{S}}$:
@@
    The optimal values satisfy
    $v^{*}(\mu,\Pi,\mathrm{K})=\mu[u^{*}].$
@@
    $v^{*}(\mu,\Pi,\mathrm{K})=\sup_{\pi\in\Pi}\inf_{\kappa\in\mathrm{K}}v^{*}(\mu,\pi,\kappa)=\inf_{\kappa\in\mathrm{K}}\sup_{\pi\in\Pi}v^{*}(\mu,\pi,\kappa).$
Theorem 2 implies that if the supremum and the infimum in the Bellman equation (3.1) interchange, then so does the sup and inf in the optimal value function. Theorem 2 is proved in Appendix A.2.
@@ 3.2.1 Sion’s Minimax Principle
In light of the exchange of the order of inf-sup in (3.1) and (3.2), a natural sufficient condition that can guarantee the assumption of Theorem 2 can be deduced from Sion’s minimax principles (Sion, 1958).
We consider the $s$-indexed family of functionals $\left\{{w_{s}^{*}:s\in S}\right\}$ defined as follows: For $d\in\mathcal{Q}$ and $p_{s}\in\mathcal{P}_{s}$
$w_{s}^{*}(d,p_{s}):=d[r(s,\cdot)]+d\otimes p_{s}[u^{*}].$ (3.5)
Clearly, $w_{s}^{*}$ is continuous, and $w_{s}^{*}(\cdot,p_{s})$ is linear in the sense of vectors in $\mathbb{R}^{A}$. Moreover, $p_{s}\in\mathcal{P}_{s}\subset\left\{{A\rightarrow\mathcal{P}(\mathcal{S})}\right\}$ can be seen as a vector-valued function $A\rightarrow\mathbb{R}^{|S|}$ which is a vector space, and $w_{s}^{*}(d,\cdot)$ is affine. Therefore, we can apply Sion’s minimax theorem if $\mathcal{Q}$ and $\mathcal{P}_{s}$ satisfy appropriate convexity and compactness assumptions.
@@
We say that a set of action distributions $\mathcal{Q}\subset\mathcal{P}(\mathcal{A})$ is convex if for all $d,d^{\prime}\in\mathcal{Q}$,
$\left\{{td+(1-t)d^{\prime}:t\in[0,1]}\right\}\subset\mathcal{Q}.$
Moreover, the set of adversarial decision rules $\mathcal{P}_{s}$ is convex if for all $p_{s},p_{s}^{\prime}\in\mathcal{P}_{s}$,
$\left\{{tp_{s}+(1-t)p_{s}^{\prime}:t\in[0,1]}\right\}\subset\mathcal{P}_{s}.$
@@
Assume that $\mathcal{Q}$ and $\mathcal{P}_{s}:s\in S$ are convex in the sense in Definition 6. Moreover, assume that either $\mathcal{Q}$ or all $\mathcal{P}_{s}:s\in S$ are compact. Then, the solution $u^{*}$ of (3.1) satisfies (3.2). Hence the conclusions of Theorem 2 holds.
Proof of Corollary 2.1.
A direct application of Sion’s minimax theorem (Corollary 3.3 in Sion (1958)) to each of $\left\{{w_{s}^{*}:s\in S}\right\}$ yields
$u^{*}(s)=\sup_{d\in\mathcal{Q}}\inf_{p_{s}\in\mathcal{P}_{s}}w_{s}^{*}(d,p_{s})=\inf_{p_{s}\in\mathcal{P}_{s}}\sup_{d\in\mathcal{Q}}w_{s}^{*}(d,p_{s})$
for all $s\in S$. Therefore, $u^{*}$ satisfies (3.2). ∎
@@
The convexity and compactness assumption in Corollary 2.1 is a sufficient condition for $u^{*}$ satisfying both (3.1) and (3.2). Many S-rectangular DRMDP models of engineering systems naturally have a convex adversary under mild assumptions, c.f. the inventory model in Section 4 of Shapiro (2022).
@@ 3.2.2 The SA-Rectangular Setting
Next, we focus on the important special case where the adversary is SA-rectangular with sets of adversarial decision rules $\left\{{\mathcal{P}_{s,a}\in\mathcal{P}(\mathcal{S}):s\in S,a\in A}\right\}$.
Note that the SA-rectangular setting, in general, doesn’t satisfy Corollary 2.1 as the set $\mathcal{P}_{s}=\prod_{a\in A}\mathcal{P}_{s,a}$ can be non-convex if $\mathcal{P}_{s,a}$ is not convex. Moreover, we allow the controller to be non-convex as well; e.g. $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}$. However, the sup-inf in the Bellman equation does interchanges. A generalization of this is stated below.
@@
$q^{*}(s,a)=r(s,a)+\gamma\inf_{p_{s,a}\in\mathcal{P}_{s,a}}p_{s,a}\left[u^{*}\right]$ (3.6)
@@
Theorem 3 implies that if $\mathcal{P}_{s}=\prod_{a\in A}\mathcal{P}_{s,a}$ for all $s\in S$ are SA-rectangular and the controller is allowed to take on deterministic policies, then a DPP always holds with deterministic decision rules. We defer the proof of this theorem to Appendix A.3.
@@ 3.3 Convex Set of Controller’s Action
We have proved the validity of the check marks for Table 4. One notices the difference between Table 3 and 4, where the (1,2)th entry, i.e. history-dependent controllers with Markov adversaries, becomes a check. In this section, we prove its validity by establishing a general principle for controllers with a convex set of action distributions $\mathcal{Q}$ defined in Definition 6. We highlight the effect of convexity of $\mathcal{Q}$ by introducing the following proposition.
Proposition 2.
$v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{H}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})\leq\mu[u^{*}]$
for all $\mu\in\mathcal{P}(\mathcal{S})$.
The main message of Proposition 2 is that when the adversary works against a convex history-dependent controller, it is sufficient for it to pick only Markov adversarial decision rules to make the value smaller than the solution of the Bellman equation. The proof of this key proposition is deferred to Appendix B.3.
An immediate consequence of Proposition 2 and Theorem 1 is the following Theorem:
@@
Assume the assumptions of Proposition 2. Then, we have that $\mu[u^{*}]=v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{H}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})$ in addition to the equalities in Theorem 1.
Proof of Theorem 4.
Given Theorem 1, it remains to show that $v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{H}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})=\mu[u^{*}]$. By Theorem 1, Lemma 1, and Proposition 2, we have
$\mu[u^{*}]=v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})\leq v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{H}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})\leq\mu[u^{*}].$

### v4

@@ 3.1.1 Deterministic Controller Policies
An important special case is when $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}=\left\{{\delta_{\left\{{a}\right\}}:a\in A}\right\}$. A deterministic rule is naturally uniquely identified with a function $\pi_{t}:\mathbf{H}_{t}\rightarrow A$. We will always assume such identification and use $\pi_{t}(h_{t})$ as a delta measure or an action interchangeably.
As clarified in [79], in the S-rectangular setting with $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}$ for $\Pi$ and $\mathcal{Q}=\mathcal{P}(\mathcal{A})$ for $\Pi$, $v(\mu,\Pi,\mathrm{K})<v(\mu,\Pi,\mathrm{K})$ could happen for $\mathrm{K}=\mathrm{K}_{\mathrm{H}}^{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$. Thus, there could be no deterministic policy that is optimal for the controller that is allowed to take randomized actions. However, if one is to constrain the controller to take deterministic action, it is still instrumental to define the $q$-function and its Bellman equation as in the classical MDP settings.
@@
Define the robust Bellman equation for the $q$-function as
$q(s,a)=\inf_{p_{s}\in\mathcal{P}_{s}}E_{\delta_{a},p_{s}}\left[r(s,A_{0})+\gamma\max_{b\in A}q(X_{1},b)\right].$ (3.3)
It is important to note that for each $s$, the infimum chooses a $p_{s}=\left\{{p_{s,a}:a\in A}\right\}$ for each $a\in A$; i.e., such a choice could be different across $a$. Effectively, for different controller’s action $a\in A$,
$q(s,a)=r(s,a)+\inf_{\psi\in\left\{{p_{s,a}:p_{s}\in\mathcal{P}_{s}}\right\}}E_{\psi}\left[\gamma\max_{b\in A}q(X_{1},b)\right],$ (3.4)
where $X_{1}\sim\psi$ under $E_{\psi}$. In particular, $\left\{{p_{s,a}:p_{s}\in\mathcal{P}_{s}}\right\}$ is the marginalization (see (2.11)) of $\mathcal{P}_{s}$.
With the robust Bellman equation for the $q$-function defined, we explore the connection of its solution with that of the Bellman equation of the value function (3.1). Let $u^{*}$ denote the unique solution of (3.1). Then, we define the optimal $q$-function as follows: for each $(s,a)\in S\times A$
$q^{*}(s,a)=\inf_{p_{s}\in\mathcal{P}_{s}}E_{\delta_{a},p_{s}}\left[r(s,A_{0})+\gamma u^{*}(X_{1})\right]=r(s,a)+\inf_{\psi\in\left\{{p_{s,a}:p_{s}\in\mathcal{P}_{s}}\right\}}E_{\psi}\left[\gamma u^{*}(X_{1})\right].$
The following corollary links $q^{*}$ and $u^{*}$ to the solution of (3.4), and hence establishes an alternative equivalent formulation of the DPP in the case of a deterministic controller.
Corollary $1.$
$q^{*}$ is the unique solution to (3.4). Moreover, if $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}$, then $u^{*}(\cdot)=\max_{a\in A}q^{*}(\cdot,a)$ and Theorem 1 holds.
Corollary 1.1 verifies the six check marks in Table 4. Subsequently, we will examine and establish the optimality attributed to the greedy policies derived from the optimal $q^{*}$-function in Remark 4. Moreover, we will provide counterexamples in Section 5.3, substantiating the red crosses within Table 4. The proof of Corollary 1.1 is provided in Appendix B.2.
@@ 3.2 The Inf-Sup Equation and the Minimax Theorem
In this section, we revisit the setting where $\mathcal{Q}\subset\mathcal{P}(\mathcal{A})$ is an arbitrary subset of action distributions. Our focus now shifts to establishing the DPP under an additional assumption that the solution $u^{*}$ of (3.1) also satisfies (3.2).
Besides establishing the DPP within Table 2, we explore a related fundamental question: At this level of generality, does the optimal maxmin control value defined in (2.14) always satisfy a minimax theorem? In other words, can the sup and inf operations over the respective entire policy classes in (2.14) always be interchanged without altering the value? We provide an affirmative answer in Theorem 2.
@@
Let $u^{*}$ be the solution of the robust Bellman equation (3.1). Assume $u^{*}$ further satisfies the inf-sup equation (3.2). Then, the following properties holds for all $\mu\in\mathcal{P}(\mathcal{S})$ and any combination of $\mathrm{K}=\mathrm{K}_{\mathrm{H}}^{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ and $\Pi=\Pi_{\mathrm{H}},\Pi_{\mathrm{M}},\Pi_{\mathrm{S}}$:
@@
    The optimal values satisfy $v(\mu,\Pi,\mathrm{K})=E_{\mu}[u^{*}(X_{0})]$.
@@
    $v(\mu,\Pi,\mathrm{K})=\sup_{\pi\in\Pi}\inf_{\kappa\in\mathrm{K}}v(\mu,\pi,\kappa)=\inf_{\kappa\in\mathrm{K}}\sup_{\pi\in\Pi}v(\mu,\pi,\kappa).$
We emphasize that, in particular, Theorem 2 implies that if the supremum and the infimum in the Bellman equation (3.1) interchange, then so do the sup and inf in the optimal value function. Theorem 2 is proved in Appendix B.3. The proof is constructive. Intuitively, the sup-inf Bellman equation (3.1) is used to identify a candidate optimal time-homogeneous control policy $\pi^{*}$, while the inf-sup equation (3.2) yields a candidate worst-case time-homogeneous adversary strategy $\kappa^{*}$. Theorem 2 then follows from the observation that if both equations hold, $\pi^{*}$ attains the maximum of $\pi\mapsto\inf_{\kappa\in\mathrm{K}}v(\mu,\pi,\kappa)$, and $\kappa^{*}$ attains the minimum of $\kappa\mapsto\sup_{\pi\in\Pi}v(\mu,\pi,\kappa)$, for any combination of policy classes $\mathrm{K}\in\{\mathrm{K}_{\mathrm{H}}^{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}\}$ and $\Pi\in\{\Pi_{\mathrm{H}},\Pi_{\mathrm{M}},\Pi_{\mathrm{S}}\}$.
@@ 3.2.1 Sion’s Minimax Principle
In light of the exchange of the order of inf-sup in (3.1) and (3.2), a natural sufficient condition that can guarantee the assumption of Theorem 2 can be deduced from Sion’s minimax principles [70].
We consider mappings $w_{s}^{*}:\mathcal{Q}\times\mathcal{P}_{s}\rightarrow\mathbb{R}$ for every $s\in S$ defined by
$w_{s}^{*}(\phi,p_{s})=E_{\phi,p_{s}}[r(s,A_{0})+\gamma u^{*}(X_{1})].$ (3.5)
Notice that, seeing $\phi$ as a vector and $p_{s}$ as a matrix, we have
$w_{s}^{*}(\phi,p_{s})=\sum_{a\in A}\phi(a)\left[r(s,a)+\gamma\sum_{s^{\prime}\in S}p_{s,a}(s^{\prime})u^{*}(s^{\prime})\right].$
This suggest that $w_{s}^{*}(\cdot,p_{s})$ is linear and $w_{s}^{*}(\phi,\cdot)$ is affine. Therefore, we can apply Sion’s minimax theorem if $\mathcal{Q}$ and $\mathcal{P}_{s}$ satisfy appropriate convexity and compactness assumptions.
@@
We say that a set of action distributions $\mathcal{Q}\subset\mathcal{P}(\mathcal{A})$ is convex if for all $\phi,\phi^{\prime}\in\mathcal{Q}$, $\left\{{t\phi+(1-t)\phi^{\prime}:t\in[0,1]}\right\}\subset\mathcal{Q}.$ Moreover, the set of adversarial decision rules $\mathcal{P}_{s}$ is convex if for all $p_{s},p_{s}^{\prime}\in\mathcal{P}_{s}$, $\left\{{tp_{s}+(1-t)p_{s}^{\prime}:t\in[0,1]}\right\}\subset\mathcal{P}_{s}.$
@@
Assume that $\mathcal{Q}$ and $\mathcal{P}_{s}:s\in S$ are convex in the sense in Definition 6. Moreover, assume that either $\mathcal{Q}$ or all $\mathcal{P}_{s}:s\in S$ are compact. Then, the solution $u^{*}$ of (3.1) satisfies (3.2). Hence, the conclusions of Theorem 2 hold.
Proof.
Proof of Corollary 2.1 A direct application of Sion’s minimax theorem (Corollary 3.3 in [70]) to each of $\left\{{w_{s}^{*}:s\in S}\right\}$ yields $u^{*}(s)=\sup_{\phi\in\mathcal{Q}}\inf_{p_{s}\in\mathcal{P}_{s}}w_{s}^{*}(d,p_{s})=\inf_{p_{s}\in\mathcal{P}_{s}}\sup_{\phi\in\mathcal{Q}}w_{s}^{*}(d,p_{s})$ for all $s\in S$. Therefore, $u^{*}$ satisfies (3.2). ∎
@@
The convexity and compactness assumption in Corollary 2.1 is a sufficient condition for $u^{*}$ satisfying both (3.1) and (3.2). Many S-rectangular RMDP models of engineering systems naturally have a convex adversary under mild assumptions, c.f. the inventory model in Section 4 of Shapiro, [60].
@@ 3.2.2 The SA-Rectangular Setting
Next, we focus on the important special case where the adversary is SA-rectangular with sets of adversarial decision rules $\left\{{\mathcal{P}_{s,a}\in\mathcal{P}(\mathcal{S}):s\in S,a\in A}\right\}$. Note that the SA-rectangular setting, in general, doesn’t satisfy Corollary 2.1 as the set $\mathcal{P}_{s}=\bigtimes_{a\in A}\mathcal{P}_{s,a}$ can be non-convex if $\mathcal{P}_{s,a}$ is not convex. Moreover, we allow the controller to be non-convex as well; e.g. $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}$. However, the sup-inf in the Bellman equation does interchange. A generalization of this is stated below.
@@
$q^{*}(s,a)=r(s,a)+\gamma\inf_{\psi\in\mathcal{P}_{s,a}}E_{\psi}[u^{*}(X_{1})]$ (3.6)
@@
Theorem 3 implies that if $\mathcal{P}_{s}=\bigtimes_{a\in A}\mathcal{P}_{s,a}$ for all $s\in S$ are SA-rectangular and the controller is allowed to take on deterministic policies, then a DPP always holds with deterministic decision rules. We defer the proof of this theorem to Appendix B.4.
@@ 3.3 Convex Set of Controller’s Action
We have proved the validity of the check marks for Table 4. One notices the difference between Table 3 and 4, where the (1,2)th entry, i.e. history-dependent controllers with Markov adversaries, becomes a check. In this section, we prove its validity by establishing a general principle for controllers with a convex set of action distributions $\mathcal{Q}$ defined in Definition 6. The proof is deferred to Appendix B.5.
@@
Let the set of action distributions $\mathcal{Q}$ be convex, and $u^{*}$ be the solution to (3.1). Then, we have that $E_{\mu}[u^{*}(X_{0})]=v(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})$ in addition to the equalities in Theorem 1.